OpenStudy (anonymous):

Question:(8+√5)(8-√5) Answers: 59-16√5 59 64-18√5 69

OpenStudy (angela210793):

=8^2-5=64-5=59

OpenStudy (angela210793):

(a-b)*(a+b)=a^2-b^2